For two events, M and N, P(M)=0.5, P(NIM) = 0.6, and P(N|M) = 0.4. Find P(MIN). P(MIN) = (Simplify your answer. Type an in
Answer #1

We get ,

P(M\cap N)=0.3

P(M'\cap N)=0.2

Therefore ,

P(N)=P(M\cap N)+P(M'\cap N)=0.3+0.2=0.5

P(N)=0.5

Now ,

P(M'|N)=\frac{P(M'\cap N)}{P(N)}=\frac{0.2}{0.5}=0.4

Therefore ,

P(M'|N)=0.4